In a relay race, the probability of the Galaxy team winning is 22%. In another unrelated relay race, the probability of the Komets team winning is 47%. If the possibility of a tie is not an option, the probability of the Komets losing their race and the Galaxy winning theirs is ?%
Answer: There is 12% chance that Komets losing their race and the Galaxy winning theirs.
Step-by-step explanation:
Since we have given that
Probability that Galaxy team winning = 22%
Probability that Komets team winning = 47%
Probability that Komets team losing = 100-47=53%
We need to find the probability of the Komets losing and Galaxy winning their game.
So, Probability that Komets losing and Galaxy winning their game is given by
\(\dfrac{22}{100}\times \dfrac{53}{100}\\\\=\dfrac{1166}{10000}\\\\=0.1166\\\\=0.12=12\%\)
Hence, there is 12% chance that Komets losing their race and the Galaxy winning theirs.